OpenStudy (anonymous):

Imagine that an object of mass m is initially moving in the x direction through a viscous fluid at a speed v. If the only force acting on this object is a viscous drag force whose magnitude is given by F=bv, prove that the object's x-velocity is given by vx(t)=ve^-qt and determine how the constant q depends on b and m. (this problem requires some calculus involving exponentials and logarithms)

OpenStudy (anonymous):

Hint: This is actually a motion-from-force problem. Divide both sides of the x component of Newton's second law by v, and take the indefinite integral of both sides

OpenStudy (paxpolaris):

\[F=bv\]\[\implies{ma}=bv\]\[\implies{ma \over v}=b\]

OpenStudy (paxpolaris):

\[\int\limits {ma \over v} dt=bt\]

OpenStudy (anonymous):

I think PaxPolaris can help you out with this question better than i can. But thanks for thinking i can help. Im really good at algebra or math so if you need help with that. than give me a call :) good luck

OpenStudy (paxpolaris):

dv/dt = a so \(dv = a\ dt\)

OpenStudy (paxpolaris):

\[\large \implies\int\limits \frac mvdv=bt\]

OpenStudy (anonymous):

So, b/m is equivalent to q

OpenStudy (anonymous):

I got v=e^(bt/m)

OpenStudy (paxpolaris):

\[m \ln v = bt + C\] -b/m is equivalent to q

OpenStudy (anonymous):

C=Vinitial

OpenStudy (paxpolaris):

\[\large v_x(t)=v_{initial} \cdot e^{bt/m} ...\] \[v_{initial}=e^{C/m}\] ... but that's not relevant to the question. \[-q= b/m \\ \implies q=-b/m\]

OpenStudy (anonymous):

Awesome, that makes so much more sense now. Thank you for all your help!